OpenStudy (anonymous):

Solve exactly. log(base 6)(x+1) + log(base 6)(x+2) = 1 x= ?

OpenStudy (anonymous):

I though it would work out to (x+1) = 1-(x+2)

OpenStudy (anonymous):

based on what you had told me earlier

OpenStudy (anonymous):

no this is a different situation.

OpenStudy (anonymous):

You need to use properties of logarithms in this one.

OpenStudy (dumbcow):

no first combine into 1 log

OpenStudy (anonymous):

When you have a sum of logs, you can make it into the log of a product: \[log_ba + log_bc = log_b(a\cdot c)\]

OpenStudy (dumbcow):

log_6(x+1)(x+2) = 1 (x+1)(x+2) = 6 ...

OpenStudy (anonymous):

x^2+3x+2=6

OpenStudy (anonymous):

Right, then solve the quadratic.

OpenStudy (anonymous):

subtract 6 and equal to 0?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

x^2+3x-4=0 so (x+4)(x-1)

OpenStudy (dumbcow):

perfect

OpenStudy (anonymous):

and a log cannot be negative correct? so x=1 ?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

awesome you two have been huge helps

OpenStudy (dumbcow):

np